[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun Aug 27, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Aug 28 02:35:22 UTC 2017


HSU Shift Log for Sun Aug 27 2017 22:35:15 GMT-0400 (EDT)

Shift: R

Forecaster: M Brennan

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: GOES-16 Meso sector 2 started over PTC 10 at 0015Z  
and continues through 1800 UTC.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170828/14f6de3d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list